OpenStudy (tomektomeknyc):

How to do it: If theta and beta belong to Q1, and cot theta = 12/13, sin beta = √ 2/2, tan (theta + beta) is equal to: a) - 20 b) 20 c) 25 d) - 25

OpenStudy (anonymous):

is there a typo? is cot supposed to say cos?

OpenStudy (tomektomeknyc):

it is cot

OpenStudy (anonymous):

first you'll need to use the trig identities to find what theta and beta are cot means cotangent and it = 1 / tan sin = opposite / hypotenuse cos = adjacent / hypotenuse tan = opposite / adjacent  opposite, adjacent, and hypotenuse all refer to the lengths of the sides of a right triangle

OpenStudy (tomektomeknyc):

ok, how would you finish it? I am not sure how to go on

OpenStudy (anonymous):

sorry I gotta sleep I hope I can help on another day

OpenStudy (tomektomeknyc):

sure, thank you

OpenStudy (phi):

You could use the tangent of the sum of 2 angles formula see http://www.intmath.com/analytic-trigonometry/2-sum-difference-angles.php \[ \tan(a+b) = \frac{\tan(a)+ \tan(b)}{1 - \tan(a) \tan(b) } \] you need to find the tangent of \( \theta\) and \( \beta \) to use this formula. in other words, if \[ \cot(\theta)= \frac{12}{13} \] what is \( \tan(\theta) \) ? If you know \[ \sin(\beta) = \frac{\sqrt{2}}{2} \] what is \( \tan(\beta) \) ? once you know those values, use the above formula

OpenStudy (tomektomeknyc):

Thank you for your help